Sam paid $64 for 5 similar files and 3 similar textbooks.A textbook cost $8 more than a file.What was the cost of one textbook?
1
Expert's answer
2011-10-06T13:10:29-0400
let file cost x$ then textbook cost (x+8)$ so we can write equation
5*x+3*(x+8)=64 5x+3x+24=64 8x=40 x=5$.
so one file cost 5$.
and one
textbook cost 5+8=13$